A 2.6 nC charge is at the origin and a -5.8 nC charge is at x = 1.5 cm . At what x-coordinate could you place a proton so that it would experience no net force? Would the net force be zero for an electron placed at the same position?
Answer #1

Here, for the proton to not experience any force , the net electric field at that position will be zero

the electric field will be zero left of the 2.6 nC charge

k * 2.6/x^2 = k * 5.8/(x + 1.5)^2

2.6/x^2 = 5.8/(x + 1.5)^2

solving for x

x = -3.04 cm

the position of zero electric field is -3.04 cm